clevegasta:

What mass (in grams) of ammonia will react with 2.800E3 kg of Phosphoric acid?

justjm:

Formulate the chemical equation, and balance it. Ammonia - NH3 Phosphoric acid - H3PO4 \(NH_3+H_3PO_4 \rightarrow (NH_4)_3PO_4\) Balance it: \(3NH_3+H_3PO_4 \rightarrow (NH_4)_3PO_4\) Now you can use dimensional analysis: \(2.800\ \cdot10^{3}\ kg\ \cdot\frac{1000\ g}{1\ kg}\cdot\frac{1\ mol\ H_{3}PO_{4}\ }{97.998\ g}\cdot\frac{3\ mol\ NH_{3}}{1\ mol\ H_{3}PO_{4}}=\_\_\_\)

justjm:

Oh yeah, after you get mols again, you need to convert back to grams NH4

clevegasta:

thank you :)

justjm:

No problem